# AgriGPT-VL: Agricultural Vision–Language Understanding Suite

### v2

@@ Abstract
Despite rapid advances in multimodal large language models, agricultural applications remain constrained by the scarcity of domain-tailored models, curated vision–language corpora, and rigorous evaluation. To address these challenges, we present the AgriGPT-VL Suite, a unified multimodal framework for agriculture. Our contributions are threefold. First, we introduce Agri-3M-VL, the largest vision–language corpus for agriculture to our knowledge, curated by a scalable multi-agent data generator; it comprises 1M image–caption pairs, 2M image-grounded VQA pairs, 50K expert-level VQA instances, and 15k GRPO reinforcement learning dataset. Second, we develop AgriGPT-VL, an agriculture-specialized vision–language model trained via a progressive curriculum of textual grounding, multimodal shallow/deep alignment, and GRPO refinement. This method achieves strong multimodal reasoning while preserving text-only capability. Third, we establish AgriBench-VL-4K, a compact yet challenging evaluation suite with open-ended and image-grounded questions, paired with a multi-metric evaluation and an LLM-as-a-judge framework. Experiments show that AgriGPT-VL outperforms leading general-purpose VLMs on AgriBench-VL-4K, achieving higher pairwise win rates in the LLM-as-a-judge evaluation. Meanwhile, it remains competitive on the text-only AgriBench-13K with no noticeable degradation of language ability. Ablation studies further confirm consistent gains from our alignment and GRPO refinement stages. We will open-sourced all of the resource to support reproducible research and deployment in low-resource agricultural settings.
@@ 1 Introduction
The convergence of AI with critical sectors like agriculture presents a significant opportunity to address global challenges such as food security and sustainable resource management (Swaminathan, 2001; Foley et al., 2011; Clapp, 2020). With the increasing challenges posed by climate change, resource scarcity, and population growth, intelligent agricultural decision-making is becoming indispensable (Godfray et al., 2010; Rockström et al., 2017; Fan & Rue, 2020). In recent years, multimodal large language models (MLLMs) have demonstrated remarkable progress in integrating vision and language, enabling tasks such as captioning, visual question answering (VQA), and multimodal reasoning (Yin et al., 2023; Achiam & et al., 2023; Chen & et al., 2024). While Multimodal Large Language Models (MLLMs) excel at integrating vision and language on general web data (Schuhmann et al., 2022; Du et al., 2022), they are ill-equipped for the agricultural domain. The knowledge required for tasks in crop and soil science is highly specialized and absent from standard pre-training corpora (Kamilaris & Prenafeta-Boldú, 2018; Wolfert et al., 2017). Consequently, existing MLLMs struggle with agricultural terminology, exhibit factual inaccuracies, and fail to provide reliable, context-aware support for real-world farming operations (Wu et al., 2024; Rezayi et al., 2022; Yang & et al., 2025).
Several attempts have been made to build agricultural language models, such as AgriBERT (Rezayi et al., 2022), AgriLLM (Didwania et al., 2024), AgroLLM (Samuel et al., 2025), and AgroGPT (Awais et al., 2025). These efforts show the value of domain-specific adaptation but are often constrained to text-only settings or narrow task coverage. Our earlier work, AgriGPT (Yang & et al., 2025), introduced the first agriculture-specialized LLM ecosystem with a curated instruction dataset
(Agri-342K), a retrieval-enhanced reasoning module (Tri-RAG), and a benchmark suite (AgriBench-13K). While effective for textual reasoning, AgriGPT lacked visual grounding and thus could not address multimodal agricultural tasks such as pest recognition or crop diagnosis. On the other hand, general-purpose MLLMs such as InternVL (Chen et al., 2024), Qwen-VL (Team, 2024), Gemini (Achiam & et al., 2023), and LLaVA (Liu et al., 2023b) demonstrate strong vision–language capabilities but are trained primarily on internet-scale data describing common objects, scenes, and events, which fail to capture agricultural semantics. As a result, these models suffer from hallucinations, poor transferability, and lack of reasoning ability in agriculture-specific scenarios. Related domains such as medicine developed specialized multimodal LLMs, highlighting the need for a comparable ecosystem in agriculture.
@@
  • Agri-3M-VL Dataset & Data Generator. We build a transferable, reusable multi-agent Data Generator and use it to curate Agri-3M-VL: 1M image–caption pairs, 2M high-quality image-grounded VQA pairs, 50K expert-level VQA, and 15k rewarded GRPO reinforcement learning dataset(10K VQA + 5K single-choice questions). To the best of our knowledge, this is the largest agriculture vision–language corpus to date.
  • AgriGPT-VL & Curriculum Training. Using a progressive curriculum, we train the agriculture-specialized VL model AgriGPT-VL, as shown in figure 1, which surpasses most flagship models in capability. To the best of our knowledge, this is currently the only open-source agriculture vision–language large model.
  • AgriBench-VL-4K & Evaluation Framework. We construct a comprehensive and challenging benchmark with 2,018 open-ended VQA and 1,858 image-grounded single-choice questions (two per image for cross-consistency), coupled with multi-metric evaluation and a dual evaluation framework that includes an LLM-as-a-judge for pairwise preference.
  • Reproducible Resources. We will open-sourced the datasets, models, and evaluation tools to support reproducible research and deployment in low-resource agricultural settings.
@@ 2.1 Text-Only Language Models in Agriculture
Pioneering work in agricultural AI largely focused on the language modality. Early models such as AgriBERT (Rezayi et al., 2022) adapted language model pre-training to domain-specific text corpora. Subsequent efforts, including AgriLLM (Didwania et al., 2024), AgroLLM (Samuel et al., 2025), and AgriGPT (Yang & et al., 2025), advanced this paradigm by developing large-scale instruction datasets like Agri-342K and text-only benchmarks such as AgriBench-13K (Yang & et al., 2025). For instance, Zhu et al. (Zhu et al., 2024) reviewed the progression of text-only and multimodal agricultural LLMs, highlighting the transition from domain adaptation to instruction-based fine-tuning. Moreover, Yu and Lin (Yu & Lin, 2024) proposed a framework leveraging LLMs for agricultural knowledge inference and consultation, suggesting broader utility beyond QA. While these models demonstrated strong textual understanding, their primary limitation was the absence of visual grounding, restricting their applicability to tasks that do not require image interpretation.
@@ 2.2 Emergence of Multimodal Agricultural Systems
The integration of visual data marked a critical evolution in agricultural AI. Foundational datasets like PlantVillage (Hughes & Salathé, 2015) and IP102 (Wu et al., 2019) provided large-scale image collections for specific recognition tasks, such as pest and disease identification. More recent works have begun to build multimodal models and benchmarks with broader capabilities. For instance, Agri-LLaVA (Wang et al., 2024), AgriCLIP (Nawaz et al., 2024), and LLMI-CDP (Wang et al., 2025) introduced vision–language abilities, while datasets like VL-PAW (Yu et al., 2025) and benchmarks like AgMMU (Gauba et al., 2025), AgroBench (Shinoda et al., 2025), and AgriEval (Yan et al., 2025) introduced tasks such as VQA and captioning. Other studies such as Zhu et al. (2024) provide a systematic review of the current landscape, while Yu & Lin (2024) and Arshad et al. (2025) explore concrete frameworks or empirical evaluations of VLMs in agricultural use cases. However, these multimodal resources often remain limited in scale, are restricted to narrow recognition tasks, or lack rigorous, large-scale quality control, representing disparate efforts rather than a cohesive foundation.
@@ 2.3 The Need for a Unified Vision–Language Ecosystem
The limitations of prior work highlight a clear need for a comprehensive and unified framework. While previous efforts have made valuable contributions to datasets, models, or benchmarks individually, progress has been hampered by the lack of a single ecosystem that integrates all three components at scale. To address this fragmentation, our work introduces a cohesive suite of resources. Our AgriGPT-VL dataset provides scale and quality; our AgriGPT-VL model handles complex reasoning beyond simple recognition; and our AgriBench-VL-4K benchmark enables robust, multifaceted evaluation. Together, these components form the kind of unified foundation we argue is necessary for the next generation of agricultural AI.
@@ 3 AgriGPT-VL
3.1 Dataset
Figure 2: Overview of the AgriGPT-VL training pipeline and curriculum-based model evolution.
@@ 3.1 Dataset
As shown in Figure 4, we aggregated a wide range of datasets covering pests and diseases, insects, crops, weeds, and fruits. Specifically, the PlantVillage dataset contains 54,305 images across 38 classes (Abdallah, 2019). For insect-related data, we included 6,878 images covering 166 fine-grained insect species from the Species196 dataset (He et al., 2023), and the Insect Foundation dataset with 317,128 images spanning 38,867 fine-grained insect classes (Yu, 2020), totaling 324,006 images and 39,033 classes. In the crop and weed domain, the SelectDataset provides 558,930 images over 2,958 categories (Contributors, 2021). For fruits, we incorporated Fruits-360 with 97,255 images and 206 categories (Murean & Oltean, 2018), and Fresh-Rotten Fruit with 30,357 images and 18 categories (Densu341, 2022), amounting to a combined 157,969 images and 224 classes. Altogether, these datasets cover 1,064,853 images and 42,253 fine-grained categories, nearly encompassing the full agricultural visual landscape.
@@
(3) Multi-Agent Refinement. Based on the preceding Image–Caption data, we generate about two VQA pairs per image. Quality is controlled by a loop formed by three agents—Feedback, Evaluation, and Rethinking. Feedback proposes revisions, Evaluation scores samples along dimensions such as factual consistency and image grounding, and Rethinking rewrites with self-consistency checks; samples iterate among the three until preset thresholds are met, yielding roughly 2M high-quality VQA. Separately, we use GPT-4o to verify and polish an additional 50K subset for supervised fine-tuning, and we additionally construct 15k GRPO reinforcement learning dataset for reward modeling and preference optimization.
(4) Instruction Filtering. Finally, we introduce a filter agent to discard irrelevant or hallucinated instructions. For example, generic questions unrelated to the image (e.g., “How many legs does a seven-spotted ladybug have?”) are removed, while faithful image-grounded questions (e.g., “What kind of insect is in the picture?”) are retained. After filtering and manual verification, about 50K expert instructions remain, representing the most reliable and domain-specific supervision signals. This filtering step guarantees the factual alignment of data, mitigating hallucinations and improving trustworthiness in model training.
@@
3.2 Model
This section details our training paradigm for AgriGPT-VL, which follows a progressive curriculum: textual grounding first, then vision–language alignment. We first consolidate domain knowledge and instruction style on text-only data, and then align vision and language on synthesized multimodal supervision with an easy-to-hard schedule.
Stage-1 (text-only domain grounding). Starting from Qwen2.5-VL, we conduct continual pretraining on $\approx$200K documents ($\approx$2.2B tokens) to inject agricultural terminology and background knowledge, followed by supervised instruction tuning on Agri-342K (Yang & et al., 2025). A held-out split of AgriBench-13K (Yang & et al., 2025) is used for early stopping and calibration prior to multimodal alignment.
@@ 3.2 Model
(2a) Shallow Alignment (Captioning, Vision Frozen). We start with 1M image–caption pairs, keeping both the vision encoder and LLM component fully frozen. Only the connector and adapter layers are trained. Captioning tasks help establish a stable semantic bridge between vision and language modalities.
(2b) Deep Alignment (From Coarse to Full Reasoning). Next, we train on 2M image–QA samples (two questions per image for cross-validation), covering recognition, attributes, diagnosis, and basic multi-hop reasoning. Using LoRA, we gradually unfreeze the vision encoder and LLM, enabling transition to full multimodal reasoning.
(2c) GRPO Optimization (Reward-Guided Fine-Tuning). We build an expert pool via multi-agent refinement, selecting 50K GPT-4o-polished samples for supervised fine-tuning, and 15K GRPO samples for reinforcement learning. GRPO rewards image-text consistency, internal logic, and verifiable terminology. Details are in Appendix A.2 and A.6.
Figure 6: Text-only evaluation of vision-language models on AgriBench-13K.
3.3 Benchmark-VL-4K
Figure 7: Pairwise win rate of vision-language models vs. AgriGPT-VL (LLM-judged)
AgriBench-VL-4K jointly evaluates generative and discriminative abilities with two components built from agricultural images: 2,018 open-ended question–answer pairs and 1,858 single-choice questions. Each image is paired with two single-choice questions to cross-validate predictions and reduce random guessing. The benchmark is strictly de-duplicated against all training splits and then human-reviewed to ensure fairness and reproducibility.
Construction. Open-ended questions are synthesized from structured captions of held-out images, covering recognition, symptom/mechanism analysis, management recommendations, and simple multi-step reasoning; answers are normalized for synonyms, units, and terminology. For the single-choice part, two complementary questions are generated per image. Distractors are mined from confusable taxa and co-occurring conditions to increase discriminability. To mitigate leakage, stems and option templates are designed not to overlap with training prompts.
Quality control and de-duplication. We remove near-duplicates at both the image and text levels: perceptual hashing or visual-feature similarity for images, and lexical/embedding similarity for question–answer strings. De-duplication is applied across train–evaluation as well as within the evaluation split. All remaining items undergo a two-pass human review by agriculture-literate annotators focusing on factual correctness, image-grounded evidence, and ambiguity resolution, with adjudication for disagreements.

### v3

@@ Abstract
Despite rapid advances in multimodal large language models, agricultural applications remain constrained by the scarcity of domain-tailored models, curated vision–language corpora, and rigorous evaluation. To address these challenges, we present the AgriGPT-VL Suite, a unified multimodal framework for agriculture. Our contributions are threefold. First, we introduce Agri-3M-VL, the largest vision–language corpus for agriculture to our knowledge, curated by a scalable multi-agent data generator; it comprises 1M image–caption pairs, 2M VQA (Visual Question Answering) pairs, 50K expert-level VQA, and 15K GRPO reinforcement learning dataset. Second, we develop AgriGPT-VL, an agriculture-specialized vision–language model trained via a progressive curriculum of textual grounding, multimodal shallow/deep alignment, and GRPO refinement. This method achieves strong multimodal reasoning while preserving text-only capability. Third, we establish AgriBench-VL-4K, a compact yet challenging evaluation suite with a multi-metric evaluation and an LLM-as-a-judge framework. Experiments show that AgriGPT-VL outperforms leading general-purpose VLMs on AgriBench-VL-4K, achieving higher pairwise win rates in the LLM-as-a-judge evaluation. Meanwhile, it remains competitive on the text-only AgriBench-13K with no noticeable degradation of language ability. Ablation studies further confirm consistent gains from our alignment and GRPO refinement stages.
Figure 2: Overview of the AgriGPT-VL training pipeline, showing the progression from textual tuning to shallow and deep image–text alignment, followed by GRPO refinement, together with the datasets used at each stage.
@@ 1 Introduction
The convergence of AI with critical sectors like agriculture presents a significant opportunity to address global challenges such as food security and sustainable resource management [Swaminathan2001, Foley2011, Clapp2020]. With the increasing challenges posed by climate change, resource scarcity, and population growth, intelligent agricultural decision-making is becoming indispensable [Godfray2010, Rockstrom2017, FanRue2020]. In recent years, multimodal large language models (MLLMs) have demonstrated remarkable progress in integrating vision and language, enabling tasks such as captioning, visual question answering (VQA), and multimodal reasoning [Yin2023MLLMsurvey, Achiam2023GPT4, Chen2024LVLMeval]. While Multimodal Large Language Models (MLLMs) excel at integrating vision and language on general web data [Schuhmann2022LAION5B, Du2022VLPsurvey], they are ill-equipped for the agricultural domain. The knowledge required for tasks in crop and soil science is highly specialized and absent from standard pre-training corpora [Kamilaris2018DLagri, Wolfert2017BigData]. Consequently, existing MLLMs struggle with agricultural terminology, exhibit factual inaccuracies, and fail to provide reliable, context-aware support for real-world farming operations [wu2024newagronomists, Rezayi2022AgriBERT, Yang2025AgriGPT].
Several attempts have been made to build agricultural language models, such as AgriBERT [Rezayi2022AgriBERT], AgriLLM [Didwania2024AgriLLM], AgroLLM [Samuel2025AgroLLM], and AgroGPT [Awais2025AgroGPT]. These efforts show the value of domain-specific adaptation but are often constrained to text-only settings or narrow task coverage. AgriGPT [Yang2025AgriGPT], introduced the first agriculture-specialized LLM ecosystem with a curated instruction dataset(Agri-342K), a retrieval-enhanced reasoning module (Tri-RAG), and a benchmark suite (AgriBench-13K). While effective for textual reasoning, AgriGPT lacked visual reasoning ability and thus could not address multimodal agricultural tasks such as pest recognition or crop diagnosis. On the other hand, general-purpose MLLMs such as InternVL [Chen2024InternVL], Qwen-VL [QwenTeam2024QwenVL], Gemini [Achiam2023GPT4], and LLaVA [Liu2023LLaVA] demonstrate strong vision–language capabilities but are trained primarily on internet-scale data describing common objects, scenes, and events, which fail to capture agricultural semantics. As a result, these models suffer from hallucinations, poor transferability, and lack of reasoning ability in agriculture-specific scenarios. Related domains such as medicine developed specialized multimodal LLMs, highlighting the need for a comparable ecosystem in agriculture.
@@
  • Agri-3M-VL Dataset & Data Generator. We build a transferable, reusable multi-agent Data Generator and use it to curate Agri-3M-VL: 1M image–caption pairs, 2M high-quality VQA pairs, 50K expert-level VQA, and 15k rewarded GRPO reinforcement learning dataset.
    To the best of our knowledge, this is the largest agriculture vision–language corpus to date.
  • AgriGPT-VL & Curriculum Training. Using a progressive curriculum, we train the agriculture-specialized VL model AgriGPT-VL, as shown in figure 1, which surpasses most flagship models in capability.
  • AgriBench-VL-4K & Evaluation Framework. We construct a comprehensive and challenging benchmark with 2,018 open-ended VQA and 1,858 image-grounded single-choice questions (two per image for cross-consistency).
  • Open Resources. All resources will be released as open-source to ensure full reproducibility and enable deployment in low-resource agricultural scenarios.
@@ 2.1 Text-Only Language Models in Agriculture
Pioneering work in agricultural AI largely focused on the language modality. Early models such as AgriBERT [Rezayi2022AgriBERT] adapted language model pre-training to domain-specific text corpora. Subsequent efforts, including AgriLLM [Didwania2024AgriLLM], AgroLLM [Samuel2025AgroLLM], and AgriGPT [Yang2025AgriGPT], advanced this paradigm by developing large-scale instruction datasets like Agri-342K and text-only benchmarks such as AgriBench-13K [Yang2025AgriGPT]. For instance, Zhu et al. [zhu2024harnessing] reviewed the progression of text-only and multimodal agricultural LLMs, highlighting the transition from domain adaptation to instruction-based fine-tuning. Moreover, Yu and Lin [yu2024framework] proposed a framework leveraging LLMs for agricultural knowledge inference and consultation, suggesting broader utility beyond QA. While these models demonstrated strong textual understanding, their primary limitation was the absence of visual reasoning ability, restricting their applicability to tasks that do not require visual interpretation.
@@ 2.2 Emergence of Multimodal Agricultural Systems
The integration of visual data marked a critical evolution in agricultural AI. Foundational datasets like PlantVillage [Hughes2015PlantVillage] and IP102 [wu2019ip102] provided large-scale image collections for specific recognition tasks, such as pest and disease identification. More recent works have begun to build multimodal models and benchmarks with broader capabilities. For instance, Agri-LLaVA [wang2024agri], AgriCLIP [nawaz2024agriclip], and LLMI-CDP [wang2025large] introduced vision–language abilities, while datasets like VL-PAW [Yu2025VL-PAW] and benchmarks like AgMMU [gauba2025agmmu], AgroBench [shinoda2025agrobench], and AgriEval [yan2025agrieval] introduced tasks such as VQA and captioning. Other studies such as zhu2024harnessing provide a systematic review of the current landscape, while yu2024framework and arshad2025leveraging explore concrete frameworks or empirical evaluations of VLMs in agricultural use cases. However, these multimodal resources often remain limited in scale, are restricted to narrow recognition tasks, or lack rigorous, large-scale quality control, representing disparate efforts rather than a cohesive foundation.
@@ 2.3 The Need for a Unified Vision–Language Ecosystem
The limitations of prior work highlight a clear need for a comprehensive and unified framework. While previous efforts have made valuable contributions to datasets, models, or benchmarks individually, progress has been hampered by the lack of a single ecosystem that integrates all three components at scale. To address this fragmentation, our work introduces a cohesive suite of resources. Our Agri-3M-VL dataset provides scale and quality; our AgriGPT-VL model handles complex reasoning beyond simple recognition; and our AgriBench-VL-4K benchmark enables robust, multifaceted evaluation. Together, these components form the kind of unified foundation we argue is necessary for the next generation of agricultural AI.
@@ 3 AgriGPT-VL
3.1 Agri-3M-VL Dataset
@@
As shown in Figure 4, we aggregated a wide range of datasets covering pests and diseases, insects, crops, weeds, and fruits. Specifically, the PlantVillage dataset contains 54,305 images across 38 classes [plantvillage2019]. For insect-related data, we included 6,878 images covering 166 fine-grained insect species from the Species196 dataset [he2023species196], and the Insect Foundation dataset with 317,128 images spanning 38,867 fine-grained insect classes [insectfoundation2020], totaling 324,006 images and 39,033 classes. In the crop and weed domain, the SelectDataset provides 558,930 images over 2,958 categories [selectdataset2021]. For fruits, we incorporated Fruits-360 with 97,255 images and 206 categories [fruits3602018], and Fresh-Rotten Fruit with 30,357 images and 18 categories [freshrotten2022], amounting to a combined 157,969 images and 224 classes. Altogether, these datasets cover 1,064,853 images and 42,253 fine-grained categories, nearly encompassing the full agricultural visual landscape.
@@
(3) Multi-Agent Refinement. To further construct high-quality VQA data on top of the Image–Caption corpus, we design a protocol-guided multi-agent refinement architecture and adopt Qwen2.5-72B as the core execution model to balance efficiency and computational cost. The architecture consists of three expert agents—Feedback, Evaluation, and Rethinking—which collaborate through a structured “generate $\rightarrow$ assess $\rightarrow$ revise” loop. The Feedback agent generates an initial question–answer draft based on the image and its caption, with protocol constraints requiring that the content reference only visible entities and attributes, avoid out-of-image inferences, and maintain clear visual evidence. The Evaluation agent agent performs structured quality assessment along correctness, clarity, and completeness (each scored within $[0,1]$) and also produces concise diagnostic feedback identifying inconsistencies, ambiguity, or insufficient grounding; any score below the unified threshold of $0.85$ triggers revision. The Rethinking agent revises the question and answer according to this feedback and performs a self-consistency check by independently generating the answer three times, requiring an agreement of at least $0.85$; otherwise, further revision is initiated. This protocolized refinement loop typically converges in approximately $2.3$ rounds, and we retain only samples that satisfy all quality requirements, yielding about 22M high-quality VQA pairs. We further select 5050K samples for multi-round verification and polishing using GPT-4o to construct a high-quality supervised fine-tuning subset, and additionally create a 1515K GRPO preference dataset, also with GPT-4o, for reinforcement learning. Empirically, about $8\%$ of initial drafts are filtered due to insufficient correctness or grounding, and manual inspection of a 55K subset confirms that all three quality dimensions consistently exceed $0.86$, demonstrating the effectiveness of the multi-agent refinement architecture in reducing hallucination and improving data reliability.
(4) Instruction Filtering. Finally, we introduce a Filter agent to automatically identify and discard instructions that are irrelevant to the image or potentially hallucinated. The agent evaluates each instruction along three protocol-defined dimensions—correctness, image-dependence, and grounding validity—and marks a sample for removal whenever any dimension falls below the required threshold. As a result, generic questions unrelated to the image (e.g., “How many legs does a seven-spotted ladybug have?”) are correctly filtered out, while questions that genuinely rely on visual evidence (e.g., “What kind of insect is in the picture?”) are retained. Quantitatively, about $0.6\%$ of the initial instructions are accurately identified as irrelevant or factually inconsistent; in a human audit of $300$ randomly sampled instructions, none of the filtered data exhibited such issues. This filtering step further strengthens factual alignment, suppresses hallucination propagation, and improves the overall trustworthiness of the training data.
@@
Figure 6: Pairwise win rate of vision-language models vs. AgriGPT-VL (Qwen2.5-VL-72B-as-a-judger)
3.2 AgriGPT-VL Model Training
As shown in figure 2. This section details our training paradigm for AgriGPT-VL, which follows a progressive curriculum: textual grounding first, then vision–language alignment. We first consolidate domain knowledge and instruction style on text-only data, and then align vision and language on synthesized multimodal supervision with an easy-to-hard schedule.
Stage-1 (Text-only). Starting from Qwen2.5-VL, we conduct continual pretraining on about 200K documents ($\approx$ 2.2B tokens) to inject agricultural terminology and background knowledge, followed by supervised instruction tuning on Agri-342K [Yang2025AgriGPT]. A held-out split of AgriBench-13K [Yang2025AgriGPT] is used for early stopping and calibration prior to multimodal alignment.
@@ 3.2 AgriGPT-VL Model Training
(2a) Shallow Alignment. We start with 1M image–caption pairs, keeping both the vision encoder and LLM component fully frozen. Only the connector and adapter layers are trained. Captioning tasks help establish a stable semantic bridge between vision and language modalities.
(2b) Deep Alignment (From Coarse to Full Training). Next, we train on 2M image–QA samples (two questions per image for cross-validation) and 50K GPT-4o-polished samples for supervised fine-tuning, covering recognition, attributes, diagnosis, and basic multi-hop reasoning. Using LoRA, we gradually unfreeze the vision encoder and LLM, enabling transition to full multimodal reasoning.
(2c) GRPO Optimization. We build 15K GRPO samples for reinforcement learning. GRPO rewards image-text consistency, internal logic, and verifiable terminology. Details are in Appendix A.2 and A.6.
3.3 AgriBench-VL-4K
Construction. To ensure the objectivity and reliability of evaluation, we construct AgriBench-VL-4K with an emphasis on data independence, procedural transparency, and rigorous quality control. The benchmark consists of two components: 2,018 open-ended question–answer pairs and 1,858 single-choice questions. All items are regenerated from held-out images that are never used during training, preventing any overlap with the distributional patterns of the training data generator. Open-ended questions are deeply rewritten from structured captions, covering recognition, symptom/mechanism analysis, management recommendations, and simple multi-step reasoning; answers are normalized for terminology, synonyms, and measurement units. For the single-choice portion, two complementary questions are designed per image, and cross-question consistency is used to reduce random guessing. Distractors are mined from confusable taxa and frequently co-occurring agricultural conditions to increase discriminability and reduce generator-style bias.
Quality control and de-duplication. To further mitigate potential data contamination, we perform strict de-duplication at both the image and text levels: perceptual hashing and visual-feature similarity are used to remove near-duplicate images, and lexical as well as embedding similarity are applied to eliminate near-duplicate question–answer strings. De-duplication is conducted both across train–evaluation splits and within the evaluation split itself, ensuring that AgriBench-VL-4K contains no residual patterns originating from the training set or the data generator. All remaining items then undergo manual review by annotators with backgrounds in agriculture and computer science, who jointly verify factual correctness, image-grounded evidence, and ambiguity resolution. In practice, approximately 7.8% of initial items (4% for open-ended QA and 11.5% for single-choice QA) are rejected due to factual inconsistencies or lack of image relevance.